In a fluorine molecule (F_2) the considerable repu
Answer #1

The reaction is:

Kp- 7.83

The Kp expression is:

K_p=\frac{P_F^2}{P_{F_2}}=7.83

Using the data you given:

\frac{P_F^2}{0.2}=7.83

Pe = 1.25

As pressure is directly proportional to concentration, we can make an ICE table:

\\ F_2\rightarrow 2F\\ 1.45...........0\\ -1.25......+2(1.25)\\ 0.2.......2.5

The dissocation percent is:

\%d=\frac{1.25}{1.45}*100\%=86.21\%